### v4

# Adiabatic Electron Transfer in the Barrierless and Marcus-Inverted Regimes

Abstract
Here it is shown that in the adiabatic limit of condensed-phase electron transfer, the onset of barrierless transition occurs at a lower driving force than predicted by the non-adiabatic Marcus formulation. Furthermore, in the adiabatic limit of the Marcus-inverted region, the standard mechanism of electron transfer becomes topologically forbidden. This behavior arises from a topological change in the mapping between the adiabatic and diabatic electronic surfaces, emerging precisely at the onset of the Marcus-inverted region. In this case, alternative mechanisms such as tunneling and non-radiative decay may dominate the rate, typically orders of magnitude slower than the rate calculated from Marcus theory.
preprint: AIP/123-QED
Electron transfer (ET) processes underpin many areas of scientific and technological relevance such as batteries [1, 2], electrocatalysis [3, 4], solar energy conversion [5], and biological redox [6]. The physics underlying ET processes in the non-adiabatic limit has been illuminated by Marcus [7, 8, 9]. The theory argues that condensed-phase electronic transitions must satisfy energy conservation and also the Franck-Condon principle that ET occurs at fixed nuclear positions. This suggests that the reaction coordinate must describe the difference in solvent polarization between the initial and final electronic states. The further assumption that the free energy of a particular electronic state is harmonic in deviations from the equilibrium polarization of that state culminates in an expression for the activation barrier [7, 8, 9].
A key assumption in many applications of Marcus theory has been that the process is non-adiabatic, which is in part because this assumption leads to elegant analytical expressions for the reaction rate based on the Fermi-Golden-Rule [10, 11]. The enigmatic Marcus inverted region, in which the reaction rate is expected to decrease at extremely high driving force as a consequence of the harmonic potentials, has also been conventionally analyzed through this non-adiabatic perspective [10].
However, the observation of strongly adiabatic reactions obeying Marcus-Hush-Chidsey rate expressions [12, 4, 13, 14, 15, 11] warrants reconsideration of the assumptions related to non-adiabadicity. In our recent work [16], we have argued that the quantum level-repulsion between the diabatic Marcus states can play a significant and underappreciated role in the ET free-energy landscape, and that the experimentally relevant reorganization $\lambda_{\text{eff}}$ may in fact be a distinct quantity from the theoretical diabatic curvature $\lambda$ described in Marcus theory. We found that the two quantities are related in the case of constant coupling by
$\lambda_{\text{eff}}=\lambda\left(1-\frac{2V}{\lambda}\right)^{2},$ (1)
or in the more general case of reaction-coordinate-dependent coupling by
$\lambda_{\rm eff}=\lambda-4V(q^{*})+\frac{4V(q_{\text{r}})^{2}}{\lambda}.$ (2)
We denote by $V$ the coupling between the electronic diabatic states when independent of the reaction coordinate $q$, and otherwise we denote by $V(q^{*})$ and $V(q_{\text{r}})$ the coupling value evaluated at the transition state and reaction minimum respectively. The activation barrier is given to a good approximation by
$E^{*}=\frac{(\lambda_{\rm eff}+\Delta G^{\circ})^{2}}{4\lambda_{\rm eff}},$ (3)
where $\Delta G^{\circ}$ is the thermodynamic bias. It is obvious that this equation reduces to the standard Marcus activation barrier in the case where the coupling is negligible
$\lambda_{\mathrm{eff}}\xrightarrow[V\to 0]{}\lambda,$ (4)
but differences can be enormous when $V$ is not much smaller than $\lambda.$ For instance, for aqueous electroreduction of carbon dioxide on a gold electrode in the presence of potassium ions, the reorganization energy has been fitted experimentally by Zhang et al. as $\lambda_{\text{eff}}=0.75$ eV [4], while Qin et al. have measured $\lambda=6.3$ eV in ab initio simulations of the same system [3]. Using these values in Eq. (1) predicts $V=2.1$ eV, which has been corroborated by our own recent ab initio calculation of the coupling value [17]. These results may be critical to quantitatively understand the kinetics of strongly adiabatic electrochemical reactions. Furthermore, in a computational study of organic semiconductors, values of $2V/\lambda>0.5$ have been reported [18], suggesting the possibility that this modification extends beyond electrochemistry to other applications of electron-transfer theory as well.
The previous analysis of Ref. [16] was confined to the adiabatic limit of the normal region, leaving open the corresponding behavior in the inverted region. This Letter addresses this case, extending our adiabatic model to address the inverted region. The theoretical finding is that there are three distinct regions of adiabatic ET characterized by categorically different behavior:
(i) $\lambda_{\text{eff}}<\Delta G^{\circ}<\lambda_{\text{eff}}$: ET is governed by the activation barrier given by Eq. (2)–(3).
(ii) $\lambda<\Delta G^{\circ}<-\lambda_{\text{eff}}$: ET is barrierless and will occur immediately.
(iii) $\Delta G^{\circ}<-\lambda$: ET in this inverted region would be forbidden according to a simple Marcus framework and can only occur via alternative routes such as nuclear tunneling or non-radiative transition [19].
We note that the idea that the coupling can suppress the ET rate in the inverted region has appeared in previous work [20, 21, 22]; yet in these treatments a non-adiabatic framework is still assumed and the coupling is treated as a perturbation. However, the large $2V/\lambda$ ratios discussed above are suggestive of a situation that should be viewed primarily in the adiabatic basis, where the Marcus transition state disappears completely.
To justify the proposed behavior in the three cases above, we proceed with the analysis presented in Ref. [16, 23], which represents the Marcus system in the basis of donor (D) and acceptor (A) diabatic states, parametrized by a solvent polarization coordinate $q$. The Hamiltonian is given by
$H(q)=\begin{pmatrix}E_{\text{D}}(q)&V(q)\\ V(q)&E_{\text{A}}(q)\end{pmatrix}=\begin{pmatrix}\lambda q^{2}&V(q)\\ V(q)&\lambda(1-q)^{2}+\Delta G^{\circ}\end{pmatrix}$, (5)
upon the diagonalization of which we obtain
$\begin{split}E_{\pm}(q)&=\frac{\lambda(2q^{2}-2q+1)+\Delta G^{\circ}}{2}\pm\\ &~~~~~~~~~~~~~~~~~~~~~~~~~\frac{1}{2}\sqrt{(\lambda(2q-1)-\Delta G^{\circ})^{2}+4V(q)^{2}}\end{split}$ (6)
where $q=0$ and $q=1$ are the minima of the donor and acceptor diabats respectively.
Figure 1 shows the diabats $E_{\text{D}}(q)$ (left dashed line) and $E_{\text{A}}(q)$ (right dashed line) from Eq. (5) as well as the adiabats $E_{-}(q)$ (lower solid line) and $E_{+}(q)$ (upper solid line) from Eq. (6) for a toy system with $\lambda=4V$, which gives $\lambda_{\text{eff}}=V$. Figure 1(a) depicts the energy landscape when $\Delta G^{\circ}=0$, corresponding to region (i). Figure 1(b) shows the corresponding landscape for $\Delta G^{\circ}=-\lambda_{\text{eff}}$, which corresponds to the transition between regions (i) and (ii). Figure 1(c) shows the landscape in region (ii), while finally Fig. 1(d) shows the landscape in region (iii).
We now justify the propositions (i)–(iii). To show (i), we recall the proof given in Ref. [16] that Eq. (3) differs from the true activation barrier obtained from $E_{-}(q)$ only on the order of correction terms $O\left(V^{n}(\Delta G^{\circ})^{3-n}/\lambda^{2}\right)$ for $n=0,1,2,3$. Since in the adiabatic limit we have $\lambda_{\text{eff}}\ll\lambda,$ Eq. (3) must hold for all $\Delta G^{\circ}<\lambda_{\text{eff}},$ and so (i) is established by Ref. [16].
Figure 1: The two level Marcus system under study, where $q$ is the non-dimensionalized reaction coordinate and $E$ is the free energy. The system is shown in (a) the normal region (b) the transition from the normal region to the barrierless region, (c) the barrierless region, and (d) the inverted region in which Marcus theory would predict no transition state.
To show (ii), we note that by the same argument as in (i), Eq. (3) must also hold when $\Delta G^{\circ}\approx\lambda_{\text{eff}},$ and so Eq. (3) predicts that the transition becomes barrierless when $\Delta G^{\circ}=-\lambda_{\text{eff}},$ which we confirm in Fig. 1(b) on the toy numerical model. As $\Delta G^{\circ}$ is further increased, we notice that no local minimum corresponding to state $D$ exists on the adiabatic ground state. The physical interpretation is that if the coupling is turned off (such as if the donor and acceptor species are sufficiently far apart), the system will reside on the donor diabat. But with the coupling turned on (such as if the species approach each other), the local minimum in the energy landscape ceases to exist, at which point the polarization coordinate will experience a force $\partial E/\partial q$ driving it to the minimum $A$ of the adiabatic surface. In this case, there will be no ET limitation in the strict sense, and the rate will be limited only by any barriers for the species’ approach, which can in general involve reaction coordinates linearly independent from the ET coordinate [17, 24].
To show (iii), which may be the most remarkable and counterintuitive proposition of this Letter, we note that it is not guaranteed that when the coupling is turned, that the electron will end up on the lower energy adiabat, even in the strong coupling limit. If the reactant coordinate is not near the diabatic crossing when the coupling is turned on, it will remain on the adiabat that most closely resembles its initial diabat regardless of whether it is the higher or lower energy one. For the same reason that ET in Marcus theory is assumed to be an isoenergetic process satisfying the Franck Condon principle [7], the system will remain trapped on the higher adiabat and no reaction will occur, with the exception of rare nuclear tunneling events [25, 10]. In this system, the criteria for the higher-energy adiabat coinciding with the $D$ diabat is precisely the criteria of being in the inverted region of the diabatic Marcus picture, because this is the point at which the mapping between the diabats and adiabats is altered.
At this point, it should be noted that the separation between regions (ii) and (iii) is expected to be exact only in the $T\rightarrow 0$ limit. When the temperature is finite, the probability of the system becoming trapped on the higher energy adiabat can be approximated by the Landau-Zener solution [10, 26]. The probability of an ET at a crossing event is given by
$P_{\text{ET}}=\text{exp}\left(-\frac{2\pi V^{2}}{\hbar\dot{q}\,|\Delta F|}\right),$ (7)
where $\dot{q}=\partial q/\partial t$ is related to the temperature and $\Delta F=(\partial E_{\text{A}}/\partial q-\partial E_{\text{D}}/\partial q)|_{q=q^{*}}.$ Indeed, the latter can easily be evaluated from $E_{\text{D}}(q)$ and $E_{\text{A}}(q)$ given in Eq. (5) by substituting in the coordinate of the diabatic crossing $q^{*}=\frac{1}{2}(1+\Delta G^{\circ}/\lambda),$ a standard result in Marcus theory. Interestingly, we obtain
$|\Delta F|=2|\Delta G^{\circ}|\approx 2\lambda,$ (8)
which shows how the probability of decays exponentially with the second order mixing parameter $V^{2}/\lambda,$ the same parameter which was found in Ref. [16] to be an important omission in standard Marcus theory. That is,
$P_{\text{ET}}\approx\text{exp}[-\alpha(V^{2}/\lambda)],~~\text{with}~\alpha=\pi/(\hbar\dot{q}).$ (9)
The reaction rate in the adiabatic inverted region (iii) will therefore be
$k_{\text{ET}}\approx\nu\exp\!\left\{-\frac{\alpha V^{2}}{\lambda}-\frac{(\lambda+\Delta G^{\circ})^{2}}{4\lambda k_{\mathrm{B}}T}\right\}$ (10)
where $\nu$ is the classical attempt frequency. The adiabatic limit implies $\alpha V^{2}/\lambda\gg 1$, and so $k_{\text{ET}}\rightarrow 0$ regardless of the value of the Marcus activation factor. These results suggest that in this limit, quantum mechanical tunneling mechanism or radiationless decay processes may dominate the rate, and these are often expected to be comparatively small [19]. A further implication of these results to liquid systems is that the coupling often depends on the distance between donor and acceptor—so if a reaction is suppressed due to high coupling at small distances, it may be more likely to occur at farther distances.
The framework presented here highlights the importance of continuing to question the standard interpretation of the Marcus inverted region [23, 27]. As Marcus recalled in his Nobel lecture, experimental evidence for the inverted-region behavior was not obtained until "almost 25 years after it was predicted" but in that interim "experimentally in some reactions 100% formation of an electronically excited state of a reaction product has been observed" [28]. The possibility that the general difficulty observing the Marcus inverted region could be due to the adiabaticity of many reactions should be considered, whether results corresponding to the classically forbidden region or the region of extended barrierless transition ($\lambda<\Delta G^{\circ}<-\lambda_{\text{eff}}$) might explain a given experiment. Tunable Hamiltonian model systems, where both the driving force and electronic coupling can be systematically varied, may provide a route toward direct experimental testing of these predictions [29, 30, 31, 32, 30].
Acknowledgments
E.A. is grateful to Martin Z. Bazant and Troy Van Voorhis for valuable support, and to Abraham Nitzan and Junghyun Yoon for useful discussions.